# Dimensional crossover of class D real-space topological invariants

### v1

@@ V Conclusions and outlook
We have established a framework based on the spectral localizer, a real-space numeric K-theory approach, to characterize topological transitions due to spatial dimension changes. In particular, we introduced local topological markers for systems in class D, with particle-hole symmetry and broken time-reversal and chiral symmetry, and applied them to Shiba lattices as experimentally-relevant examples. We considered a 2D circular Shiba island and deformed it into a 1D chain by reducing its size along one of its dimensions. By computing the local markers across the dimension change, we found that the transition from 2D topology to 1D topology occurs via a reduction in the number of sites with Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|>0$ together with a decrease in the localizer gap protecting it. At a critical thickness, we found the emergence of nontrivial 1D topology, marked by $z_{2,\bm{\lambda}}=-1$, protected by a nonzero 1D localizer gap whose size increases for the 1D chain geometry. We tested the robustness of our results against disorder in the chemical potential. We found that the 1D localizer gap, and thus the 1D topology, is more robust than its 2D counterpart.
@@
The framework introduced can be readily generalized to other symmetry classes. For example, the spectral localizer, Eq. (3), can be defined for a 3D topological insulator and a 2D quantum spin Hall system. The corresponding topological invariants would be a Chern number, analogous to Eq. (6), and the spin Chern number. More generally, the framework could help guide the design of topological materials for device applications. One of the strengths of the spectral localizer framework is its numerical efficiency, which enables it to be applied directly to experimental systems without developing a low-energy approximation [66]. As such, our results may be useful for the design and analysis of topological superconducting nanowires [67].
@@ Acknowledgements.
We thank Annica M Black-Schaffer and Arnob Kumar Ghosh for insightful discussions. A.C. acknowledges support from the Laboratory Directed Research and Development program at Sandia National Laboratories. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology &\&& Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. Research by T.L. was supported in part by the National Science Foundation, award DMS-2349959. Research by T.L. was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-25-1-0052. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government.
Appendix A Appendix A
In this appendix, we consider the case of a Shiba island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into a trivial chain, with $z_{2,\bm{\lambda}}=1$. The topological dimensional crossover occurs for the three-atoms-tall island, analogous to the cases discussed in the main text, but in this case, the system is trivial.
Figure 9: (a)-(e): Arrangement of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, (orange squares) and 2D localizer gap $\text{min}\{\left|\Sigma(L^{(2D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters used are $N=20$, $\mu=0$, $\alpha=0.8$, $\Delta=1.2$, and $JS=2$. The spectral localizer calculations use $\kappa=0.05(t/l)$.

### v2

@@ III Discussion
We have established a real-space approach based on the spectral localizer framework to characterize topological transitions due to spatial dimension changes. In particular, we have leveraged local topological markers for systems in class D, with particle-hole symmetry and broken time-reversal and chiral symmetry, to study Shiba lattices in experimentally relevant examples. We considered a 2D circular Shiba island and deformed it into a 1D chain by reducing its size along one of its dimensions. By computing the local markers across the dimension change, we found that the transition from 2D topology to 1D topology occurs via a reduction in the number of sites with Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|>0$ together with a decrease in the 2D localizer gap protecting the Chern number. At a critical thickness, we found the emergence of nontrivial 1D topology, marked by $z_{2,x}=-1$, protected by a nonzero 1D localizer gap whose size increases for the 1D chain geometry. We tested the robustness of our results against disorder in the chemical potential of the magnetic adatoms, and numerically observed that the 1D localizer gap, and thus the 1D topology, is more robust than its 2D counterpart.
There are other methods to produce local markers of $K$-theory, including some that are valid for interacting systems [75, 76]. These are, however limited to the $\mathbb{Z}$-valued invariants so cannot be used to track the same Altland-Zirnbauer symmetry class across adjacent dimensions where at least one $\mathbb{Z}_{2}$-valued invariant will be needed.
@@
The approach we have introduced can be readily generalized to other symmetry classes. For example, we also considered the one-to-zero-dimensional crossover in a BDI class model, showing the generality of our approach [see Supplementary Note 1]. The spectral localizer, Eq. (3), can also be defined for a 3D topological insulator and a 2D quantum spin Hall system [77]. Additionally, one could consider the simultaneous removal of two or more dimensions, for example, to study the topological transition between 3D and 1D class AIII systems. More generally, the framework could help guide the design of topological materials for device applications. One of the strengths of the spectral localizer framework is its numerical efficiency, which enables it to be applied directly to experimental systems without developing a low-energy approximation [78]. As such, our results may be useful for the design and analysis of topological superconducting nanowires [79].
IV Methods
IV.1 Local density of states
We compute the local density of states using the Chebyshev-Bogoliubov–de Gennes method introduced in reference [64]. For completeness, we provide here a brief description. The Green’s function is defined as $G=\langle vac|\Psi^{\dagger}\hat{G}(E)\Psi|vac\rangle$, where $\hat{G}(E)=\left[E+i\epsilon-H\right]^{-1}$, $H$ the Hamiltonian in Eq. (1), $|vac\rangle$ is the vacuum, and $\Psi=(c_{1\uparrow},...,c_{N^{2}\uparrow};c_{1\downarrow},...,c_{N^{2}\downarrow};c^{\dagger}_{1\uparrow},...,c^{\dagger}_{N^{2}\uparrow};c^{\dagger}_{1\downarrow},...,c^{\dagger}_{N^{2}\downarrow})$, labeling the positions in the two-dimensional superconductor using a single index. $G^{\sigma\sigma}_{jj}$ in Eq. (2) corresponds to the matrix entry with spin $\sigma$ at site $j$. As shown in reference [64], expanding the Green’s function using Chebyshev polynomials leads to
$G_{jj}^{\sigma\sigma}(E)$ $\approx\frac{-2i}{\sqrt{1-\tilde{E}^{2}}}\sum_{m=0}^{M}a_{m}^{\sigma\sigma}(j,j)\frac{\sinh[\nu(1-\frac{m}{M})]}{\sinh(\nu)}$
$\quad\times e^{-im\arccos(\tilde{E})}$, (10)
with coefficients $a^{\sigma\sigma}_{m}(i,j)=\langle c_{i\sigma}|T_{m}(H)|c_{j\sigma}^{\dagger}\rangle/(1+\delta_{0,m})$, where $|c_{j\sigma}^{\dagger}\rangle=c_{j\sigma}^{\dagger}|vac\rangle$, and $T_{m}(x)$ are the Chebyshev polynomials of the first kind which can be computed recursively. $\tilde{E}$ corresponds to a rescaled energy, such that the spectrum lies within the interval [1,1][-1,1]. Eq. (10) becomes exact in the limit $M\rightarrow\infty$. $\nu=\epsilon M$ is a real number taken to be the product of the Green’s function broadening $\epsilon$ and the summation cutoff $M$. We consider $\epsilon=5\times 10^{-3}$ and a cutoff $M=10^{3}$, enough for the system sizes we consider.
IV.2 Spectral localizer
On the tight-binding basis of the Hamiltonian $H$, given by $\Psi=(c_{1\uparrow},...,c_{N^{2}\uparrow};c_{1\downarrow},...,c_{N^{2}\downarrow};c^{\dagger}_{1\uparrow},...,c^{\dagger}_{N^{2}\uparrow};c^{\dagger}_{1\downarrow},...,c^{\dagger}_{N^{2}\downarrow})$, we label the positions in the system using a single index, $j=1,...,N^{2}$. Then, the position operators $X,Y$ used in the spectral localizer are diagonal matrices with the $x$ ($y$) component of the $j$-th lattice site at the entry $X_{jj}$ ($Y_{jj}$). The spin and particle-hole spaces are labeled with an additional index at the same position. The spectral localizer is diagonalized using eigenvalue solvers to obtain its spectrum and gap. The two-dimensional topological marker, $C^{L}_{\bm{\lambda}}$, is given by the signature of the spectral localizer $L^{(2D)}_{\bm{\lambda}}$. When the full spectrum is not needed, we speed up its calculation using the LDLT decomposition of the spectral localizer. To compute the sign of the determinant needed for the 1D topological marker, $z_{2,x}$, we perform a LU factorization of $L^{(1D)}_{\bm{\lambda}}$.
Data availability
The data that support the findings of this work are available from the corresponding author on reasonable request.
Code availability
The codes that support the findings in this study are available from the corresponding author upon reasonable request.
Author contributions
All authors conceived the research project. M.R-V. performed the numerical calculations, with input from A.C. and T.A.L. All authors contributed to discussions and interpretation of the results. M.R-V. and A.C. wrote the manuscript with the comments from T.A.L.
@@ Acknowledgements.
We thank Annica M Black-Schaffer and Arnob Kumar Ghosh for insightful discussions. A.C. acknowledges support from the Laboratory Directed Research and Development program at Sandia National Laboratories. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology &\& Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. Research by T.L. was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-25-1-0052. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government.
Competing interests
The authors declare no competing interests.